# Covariance Analysis of Impulsive Streaking

### v1

@@ IV.4.2 Angular Sparsity
Figure 12: Effects of sparse angular sampling on the RR delay retrieval methods. (a) The dressing-free MD collected by an array of 16 spectrometers, from the same ionization process as Fig. 1. Angular bins are grouped into the even (orange rectangles) and odd (cyan rectangles) ones. (b) Streaking covariance $K[X,Y]$ at $\phi=\pi/3$, $k/\Delta p=0.5$, under the MRCO16 measurement scheme, see main text. Radial boundaries are the same as for Fig. 3(a). (c) Same as (b) but in the MRCO8v8 scheme where $X$ and $Y$ are detected by the even and odd spectrometers, respectively. (d-f) Comparison of the three schemes (N180 in red, MRCO16 in green, MRCO8v8 in blue) on the RR (solid) and daRR (dashed) delay retrieval methods. (d) Systematic error in the noiseless limit: of the RR method (grey solid), almost the same for the three schemes, and of the daRR method (dashed, colors corresponding to the schemes). (e) In the presence of shot noise controlled by the electron flux $E_{\mathrm{xp}}$, systematic error at $k/\Delta p=0.5$ (black dotted line in (d)). (f) The solid and dashed traces are the same as (e) except for statistical errors of the RR and daRR methods, respectively. The dot-dashed traces are asymptotic statistical errors after introducing $\kappa$ sampling noise in addition to shot noise, with colors corresponding to the schemes.
@@
When we introduce shot noise to the simulations, the three measurement schemes (N180, MRCO16, MRCO8v8) share several other similarities in terms of the delay retrieval performance. Decreasing the electron flux in each set of ROIs, $E_{\mathrm{xp}}$, from the noiseless limit, we see a region where the systematic error of RR method scale as $\propto 1/E_{\mathrm{xp}}$ and then exceeds the daRR method, as shown by the solid curves in Fig. 12(e). This behavior is exhibited under all three schemes, although the values of $E_{\mathrm{xp}}$ below which the daRR method outperforms the RR method are different. As $E_{\mathrm{xp}}$ decreases further, the systematic error of the daRR method remains insensitive to $E_{\mathrm{xp}}$ until the electron counts become excessively scarce $E_{\mathrm{xp}}\lesssim 5$. On the other hand, the statistical error of both the RR and the daRR methods follows the $1/\sqrt{N_{s}E_{\mathrm{xp}}}$ scaling, as shown by the colored traces in panel (f). Remarkably, in this asymptotic scaling region the statistical error curves for the different measurement schemes overlap each other, rather than having an offset between them. Meanwhile, in the systematic error curves of the RR method, the $\propto 1/E_{\mathrm{xp}}$ scaling regions also overlap with each other when compared at the same $E_{\mathrm{xp}}$. Note that $E_{\mathrm{xp}}$ represents the total electron counts detected in a set of angular bins, thus achieving the same $E_{\mathrm{xp}}$ requires a higher total number of electrons generated per shot when the measurement scheme is more sparse. Another common feature for measurement schemes is the value of $E_{\mathrm{xp}}/N_{Q}$ where the RR and daRR methods are equally accurate. For the $k/\Delta p=0.5$ case shown in panel (e), this cross-over is at $E_{\mathrm{xp}}/N_{Q}\approx 10$. This agreement is universal so long as $k/\Delta p<0.6$ and the RR method has not reached the noiseless limit.
Next, we consider sampling noise of $\kappa$ in addition to shot noise. Due to this sampling noise of $\kappa$, when the number of shots $N_{s}$ is fixed and $E_{\mathrm{xp}}$ is high, the statistical error of retrieved delay approaches the lower bound $\sigma[\Delta\phi]\approx\sqrt{\mathcal{C}_{\kappa}/N_{s}}$ for both the RR and daRR methods, similar as in Fig. 10(b). As shown by the dot-dashed lines in Fig. 12(f), this lower bound at $\sqrt{\mathcal{C}_{\kappa}/N_{s}}$ is roughly the same under different levels of angular sparsity. Meanwhile, the overlap of the colored curves in Fig. 12(f) indicates that $\mathcal{C}_{E}$ is also almost independent of the angular sparsity. Thus in the presence of both shot noise and sampling noise of $\kappa$, the asymptote of the statistical error $\sigma[\Delta\phi]\approx\sqrt{\mathcal{C}_{E}/(N_{s}E_{\mathrm{xp}})+\mathcal{C}_{\kappa}/N_{s}}$ is not significantly affected by the angular sparsity.
@@
V Encoding Arbitrary Periodic Signal
In this section, we extend our discussion to the situation where the emission process resulting in feature $Y$ is no longer impulsive, i.e.not necessarily much shorter than $T_{L}$. The time-scale for the emission must remain shorter than the duration of the dressing field envelope. The $Y$ feature is assumed to remain in the impulsive limit, and is used as a timing reference for $X$. For $Y$ to qualify as a timing reference for the non-impulsive feature $X$, it is implied that the two processes producing the electrons in $X$ and $Y$, if not the same process, have a fixed relative temporal relation with each other. Thus the unique stationary phase time $t_{s}$ in producing electrons in $X$ determines the relative timing of the joint process producing both features, with respect to the dressing field. Furthermore, feature $Y$ is periodic with $t_{s}$ at the dressing field period $T_{L}$, assuming the duration of $|A_{0}(t)|$ remains much longer than the joint process. Thus feature $Y$ is a function of the streaking vector $\bm{k}$ of the concurrent reference $X$ feature, $Y=Y(\bm{r};\bm{k})$, which also indicates the periodicity with streaking direction.
For this study, we will consider the encoding of a feature $\mathbb{E}\left[Y|\bm{k}\right]$ in the streaking covariance $K[X,Y]$, after averaging over machine fluctuations. Assuming periodicity over $\kappa\in[0,2\pi]$, implies that we are encoding the Fourier coefficients
$\mathcal{Y}_{m}(\bm{r};k)\equiv\int_{0}^{2\pi}\mathbb{E}\left[Y(\bm{r};\bm{k})|\bm{k}\right]e^{-im\kappa}\frac{d\kappa}{2\pi}~{},$ (19)
in to the measured electron momentum distribution (MD). Applying the law of total covariance to $K[X,Y]$ by conditioning on the streaking amplitude $k$, we observe that all the Fourier components $\mathcal{Y}_{m}$ are encoded in $K[X,Y]$, but not always with the same sensitivity:
@@ V Encoding Arbitrary Periodic Signal
$K[X,Y]$ $=\sum_{m=1}^{+\infty}\left(\langle\mu_{m}(k)\mathcal{Y}_{m}(k)\rangle+\mathrm{c.c.}\right)+C\left[\mu_{0}(k),\mathcal{Y}_{0}({k})\right]$ (20)
@@
$\mu_{m}(k)$ $\equiv\frac{1}{m!}\left(\frac{-k\partial_{+}}{\sqrt{2}}\right)^{m}{}_{0}F_{1}\left(m+1,\frac{k^{2}\nabla^{2}}{4}\right)\langle X^{0}\rangle~{},\quad m\geq 0$ (21)
@@
where $\langle X^{0}\rangle$ is still the dressing-free MD averaged over machine fluctuations, $\partial_{+}=(\frac{\partial}{\partial x}+i\frac{\partial}{\partial y})/\sqrt{2}$ is an operator on a MD, and ${}_{0}F_{1}(z,x)\equiv\sum_{n=0}^{\infty}((z-1)!x^{n})/((z+n-1)!n!)$ is the confluence hypergeometric limit function [26]. Equations (20) and (21) only require the periodicity of $\mathbb{E}\left[Y(\bm{r};\bm{k})|\bm{k}\right]$ with $\kappa$, as shown in Supplemental Sec. 4. This indicates that an arbitrary periodic signal of $\kappa$, not limited to electron yield in a momentum region in an angular streaking measurement, is encoded in the streaking covariance with the impulsive reference feature $X$ in the same way as the encoding of $\mathbb{E}\left[Y(\bm{r};\bm{k})|\bm{k}\right]$ described in Eqn. (20).
@@
In writing Eqn. (20) we have partitioned the terms into $\kappa$-alternating terms ($|m|{\geq}1$) and the $\kappa$-averaged terms ($m{=}0$) Inspecting Eqn. (20), it can be seen that the $|m|{\geq}1$ components of $Y$ are encoded only in the first term, resulting from the shot-to-shot variation of streaking direction $\kappa$. The streaking covariance $K[X,Y]$ depends linearly on these $|m|{\geq}1$ components, with sensitivity $\mu_{m}(k)$ given by Eqn. (21). In contrast, the $m{=}0$ component of $Y$ only affects the second term $C\left[\mu_{0}(k),\mathcal{Y}_{0}({k})\right]$, which indicates that $K[X,Y]$ is sensitive to the covariance of $\mathcal{Y}_{0}$ and $\mu_{0}$ arising from the variation of streaking amplitude $k$. Note that in the small streaking regime ($k<\Delta p_{X}$ of the reference feature), the leading order of the sensitivity is $\mu_{m}\approx\frac{k^{m}}{m!2^{m/2}}\partial_{+}^{m}\langle X^{0}\rangle$, where the minimal order of differentiation on $\langle X^{0}\rangle$ is $m$. Thus only the components $\mathcal{Y}_{\pm 1}$ are coupled to the gradient of the reference feature $\nabla\langle X^{0}\rangle$. This indicates that the delay retrieval methods discussed in the preceding text can also be considered as an analysis of the $|m|=1$ components.
@@
Returning to Eqn. (20), the sensitivity to the $|m|\geq 2$ components is typically lower than to the $|m|=1$ components in the small streaking regime $k<\Delta p_{X}$. To illustrate this point, we simulate the MD of an feature $Y$ that is periodic over $\kappa\in[0,2\pi]$. We consider the Auger-Meitner (AM) decay of a molecule in a superposition of two core-ionized states, prepared by core ionization by a $0.2$ fs FWHM pulse. The two intermediate cationic states are spaced by $\Delta I_{\mathrm{P}}=7\hbar\omega_{L}$, and from these two states to a common dication state, the dressing-free AM electrons are centered at $4.34$ a.u. and $4.38$ a.u. in momentum, respectively. The molecules are ionized in the presence of a $|\bm{A}|=0.1$ a.u., $cT_{L}=1.85\,\mathrm{\mu m}$ dressing field (same as Fig. 10) and the MD is measured in the $p_{z}=0$ slicing measurement scheme. Other details of the simulation of AM distribution are elaborated in Supplemental Sec. VI. The decay of this superposition of core-ionized states results in a characteristic modulation of $Y$ over $\kappa$ at a period of $72\pi/$ 7, which results from the interference between the pathways via two different intermediate states [27]. This modulation appears across a wide range of AM momenta. In fact, because $Y^{0}$ has been assumed to be angularly isotropic, the MD of AM electrons at different angular positions $\theta_{p}$ are all equivalent up to an offset in $\kappa$, i.e. $Y(\theta_{p},r;\kappa,k)=Y(0,r;\kappa-\theta_{p},k)$, so the yield is also modulated over angular position $\theta_{p}$.
@@
For the purpose of illustration, we set the $N_{P}=180$ ROIs on feature $Y$, $\{P_{p}\}$, between $p_{\mathrm{min}}{=}4.40$ a.u. and $p_{\mathrm{max}}{=}4.50$ a.u.. Across different streaking directions of the reference feature $\kappa\in[0,2\pi]$, the resultant yield in the angular bin at $\theta_{p}=0$ is shown in Fig. 13(a). The yield in these angular bins $Y_{p}$ is clearly modulated with $\kappa$ and $\theta_{p}$, as shown in panel (b). We compute the streaking covariance $K[X,Y]$, employing the same photoelectron feature $X$ that was employed in Fig. 10. This reference feature $X$ is simulated in the same dressing field ($k=0.1\,\mathrm{a.u.}\approx 0.5\Delta p_{X}$), and the ROIs $\{Q_{q}\}$ are still bounded between $p_{\mathrm{MG}}$ and $p_{\max}{=}2.15$ a.u.. As shown in Fig. 13(c), $K[X,Y]$ barely displays the $72\pi/$ 7-period modulation over either $\theta_{p}$ or $\theta_{q}$. In contrast to the amplitude ratio between the $m{=}7$ and $m{=}1$ Fourier components of $Y$ is $|\mathcal{Y}_{7}|/|\mathcal{Y}_{1}|=0.44$, the amplitude ratio between the $m{=}7$ and $m{=}1$ Fourier components of $K[X|_{\theta_{q}=0},Y]$ has reduced to $2\times 10^{-5}$. Such a significant attenuation of the $72\pi/$ 7-period modulation results from the low sensitivity $\mu_{7}$ in this case. We compute the sensitivity $\mu_{m}$ according to Eqn. (21) and then integrate over the angular bins $Q_{q}$, and as shown in panel (d), the magnitude $|\mu_{m}|$ overall decreases with the increase of $s$ in this case. The products of multiplying these calculated $\mu_{m}$ to the corresponding Fourier components $\mathcal{Y}_{m}$ yield good agreements with the Fourier coefficients of $K[X,Y]$ with respect to $\theta_{p}$, which is visualized for the $\theta_{q}=0$ slice in panel (e), corroborating Eqn. (21).
@@
We notice that in the small streaking regime, the sensitivity is proportional to the regional integral of $\partial_{+}^{m}\langle X^{0}\rangle$, i.e. the $m$-th order derivative of the reference feature. The scaling of the sensitivity over $m$ therefore depends on the streaking amplitude relative to the characteristic momentum scale of $\langle X^{0}\rangle$, denoted as $\delta p$, and the diameter of an angular bin $d_{p}$. The magnitude of the sensitivity $|\mu_{m}|$ roughly scales as $\propto(k/\max(d_{p},\delta p))^{m}$ in the small streaking regime, which is an exponential decrease with increasing order $m$ when $k<\max(d_{p},\delta p)$. This indicates that to achieve high sensitivities to the high-$m$ components, it is desirable to employ a strong dressing field and limit the size of the ROIs on the reference feature such that $k>\max(d_{p},\delta p)$.
@@
The bilinearity of covariance also leads to the independence of sensitivity $\mu_{m}$ on feature $Y$. Rather, $\mu_{m}$ is determined by the reference feature $X$ and the streaking amplitude $k$, as shown in Eqn. (21). Therefore, when detecting $Y$ as a function of $\kappa$ by analyzing the covariance with a reference feature $X$, the sensitivity to certain Fourier components $\mathcal{Y}_{m}$ can be optimized in simulation, by tuning the streaking amplitude $k$, the shape of $X^{0}$ and the position of ROIs $\{Q_{q}\}$, without detailed prior knowledge in $Y$.

### v2

@@ IV.4.2 Angular Sparsity
Figure 12: Effects of sparse angular sampling on the RR delay retrieval methods. (a) The dressing-free MD collected by an array of 16 spectrometers, from the same ionization process as Fig. 1. Angular bins are grouped into the even (orange rectangles) and odd (cyan rectangles) ones. (b) Streaking covariance $K[X,Y]$ at $\phi=\pi/3$, $k/\Delta p=0.5$, under the MRCO16 measurement scheme, see main text. Radial boundaries are the same as for Fig. 3(a). (c) Same as (b) but in the MRCO8v8 scheme where $X$ and $Y$ are detected by the even and odd spectrometers, respectively. (d-f) Comparison of the three schemes (N180 in red, MRCO16 in green, MRCO8v8 in blue) on the RR (solid) and daRR (dashed) delay retrieval methods. (d) Systematic error in the noiseless limit: of the RR method (grey solid), almost the same for the three schemes, and of the daRR method (dashed, colors corresponding to the schemes). (e) In the presence of shot noise controlled by the electron flux $E_{\mathrm{fx}}$, systematic error at $k/\Delta p=0.5$ (black dotted line in (d)). (f) The solid and dashed traces are the same as (e) except for statistical errors of the RR and daRR methods, respectively. The dot-dashed traces are asymptotic statistical errors after introducing $\kappa$ sampling noise in addition to shot noise, with colors corresponding to the schemes.
@@
When we introduce shot noise to the simulations, the three measurement schemes (N180, MRCO16, MRCO8v8) share several other similarities in terms of the delay retrieval performance. Decreasing the electron flux in each set of ROIs, $E_{\mathrm{fx}}$, from the noiseless limit, we see a region where the systematic error of RR method scale as $\propto 1/E_{\mathrm{fx}}$ and then exceeds the daRR method, as shown by the solid curves in Fig. 12(e). This behavior is exhibited under all three schemes, although the values of $E_{\mathrm{fx}}$ below which the daRR method outperforms the RR method are different. As $E_{\mathrm{fx}}$ decreases further, the systematic error of the daRR method remains insensitive to $E_{\mathrm{fx}}$ until the electron counts become excessively scarce $E_{\mathrm{fx}}\lesssim 5$. On the other hand, the statistical error of both the RR and the daRR methods follows the $1/\sqrt{N_{s}E_{\mathrm{fx}}}$ scaling, as shown by the colored traces in panel (f). Remarkably, in this asymptotic scaling region the statistical error curves for the different measurement schemes overlap each other, rather than having an offset between them. Meanwhile, in the systematic error curves of the RR method, the $\propto 1/E_{\mathrm{fx}}$ scaling regions also overlap with each other when compared at the same $E_{\mathrm{fx}}$. Note that $E_{\mathrm{fx}}$ represents the total electron counts detected in a set of angular bins, thus achieving the same $E_{\mathrm{fx}}$ requires a higher total number of electrons generated per shot when the measurement scheme is more sparse. Another common feature for measurement schemes is the value of $E_{\mathrm{fx}}/N_{Q}$ where the RR and daRR methods are equally accurate. For the $k/\Delta p=0.5$ case shown in panel (e), this cross-over is at $E_{\mathrm{fx}}/N_{Q}\approx 10$. This agreement is universal so long as $k/\Delta p<0.6$ and the RR method has not reached the noiseless limit.
Next, we consider sampling noise of $\kappa$ in addition to shot noise. Due to this sampling noise of $\kappa$, when the number of shots $N_{s}$ is fixed and $E_{\mathrm{fx}}$ is high, the statistical error of retrieved delay approaches the lower bound $\sigma[\Delta\phi]\approx\sqrt{\mathcal{C}_{\kappa}/N_{s}}$ for both the RR and daRR methods, similar as in Fig. 10(b). As shown by the dot-dashed lines in Fig. 12(f), this lower bound at $\sqrt{\mathcal{C}_{\kappa}/N_{s}}$ is roughly the same under different levels of angular sparsity. Meanwhile, the overlap of the colored curves in Fig. 12(f) indicates that $\mathcal{C}_{E}$ is also almost independent of the angular sparsity. Thus in the presence of both shot noise and sampling noise of $\kappa$, the asymptote of the statistical error $\sigma[\Delta\phi]\approx\sqrt{\mathcal{C}_{E}/(N_{s}E_{\mathrm{fx}})+\mathcal{C}_{\kappa}/N_{s}}$ is not significantly affected by the angular sparsity.
@@
V Encoding Arbitrary Signals
In this section, we extend our discussion to the situation where the emission process approaches or exceeds the period of the dressing laser field, and thus the emission is no longer impulsive. We will remain in the limit that the emission process is shorter than the duration of the dressing field envelope. For this method to work, we still require an impulsive emission process to provide a reference feature. Moreover, these two emission processes should be correlated for the instantaneous process to serve as a reference. Similar to the conventions used above, we will refer to the impulsive feature as feature $X$, and the longer timescale emission process will produce feature $Y$.
Owing to the periodic nature of the interaction of the laser field, the time-dependence of the emission is encoded in the Fourier coefficients of the measured electron momentum distribution (MD):
$\mathcal{Y}_{m}(\bm{r};k)\equiv\int_{0}^{2\pi}\mathbb{E}\left[Y(\bm{r};\bm{k})|\bm{k}\right]e^{-im\kappa}\frac{d\kappa}{2\pi}$ (19)
Here $Y(\bm{r};\bm{k})$ is the yield of electrons belonging to feature $Y$ at detector coordinate $\bm{r}$ for a streaking vector $\bm{k}$, which is similar to the conventions used above. In writing this expression, we have taken into account the machine fluctuations by using $\mathbb{E}\left[Y|\bm{k}\right]$. Applying the law of total covariance, we can write $K[X,Y]$ as
@@ V Encoding Arbitrary Signals
$K[X,Y]$ $\simeq\sum_{m=1}^{+\infty}\left(\langle\mu_{m}(k)\mathcal{Y}_{m}(k)\rangle+\mathrm{c.c.}\right)+C\left[\mu_{0}(k),\mathcal{Y}_{0}({k})\right]$ (20)
@@
$\mu_{m}(k)$ $\equiv\frac{1}{m!}\left(\frac{-k\partial_{+}}{\sqrt{2}}\right)^{m}{}_{0}F_{1}\left(m+1,\frac{k^{2}\nabla^{2}}{4}\right)\langle X^{0}\rangle,~{}m\geq 0$ (21)
@@
where $\langle X^{0}\rangle$ is the dressing-free MD averaged over machine fluctuations, $\partial_{+}=(\frac{\partial}{\partial x}+i\frac{\partial}{\partial y})/\sqrt{2}$ is an operator that acts on a MD, and ${}_{0}F_{1}(z,x)\equiv\sum_{n=0}^{\infty}((z-1)!x^{n})/((z+n-1)!n!)$ is the confluence hypergeometric limit function [28]. It is clear from Eqn. (21) that all Fourier components, $\mathcal{Y}_{m}$, are encoded in $K[X,Y]$, but not always with the same sensitivity. This indicates a signal which is periodic with $\kappa$ is encoded in the streaking covariance with the impulsive reference feature $X$ in the same way as the encoding of $\mathbb{E}\left[Y(\bm{r};\bm{k})|\bm{k}\right]$ described in Eqn. (20). This encoding in all Fourier components is a general property of a periodic signal.
@@
In applying the law of total covariance to partition $K[X,Y]$, the conditional variable was the streaking amplitude $k$. In writing Eqn. (20) we have partitioned the terms into two terms, one that results from the shot-to-shot variations of $\kappa$ (the sum over $m{\neq}0$) and the other that results from variations in $k$. The streaking covariance $K[X,Y]$ depends linearly on the $|m|{\geq}1$ components, with sensitivity $\mu_{m}(k)$ given by Eqn. (21). In contrast, the $m{=}0$ component of $Y$ only affects the second term $C\left[\mu_{0}(k),\mathcal{Y}_{0}({k})\right]$, which indicates that $K[X,Y]$ is sensitive to the covariance of $\mathcal{Y}_{0}$ and $\mu_{0}$ arising from the variation of the streaking amplitude $k$. Note that in the small streaking regime ($k<\Delta p_{X}$ of the reference feature), the leading order of the sensitivity is $\mu_{m}\approx\frac{k^{m}}{m!2^{m/2}}\partial_{+}^{m}\langle X^{0}\rangle$, where the minimal order of differentiation on $\langle X^{0}\rangle$ is $m$. Thus only the components $\mathcal{Y}_{\pm 1}$ are coupled to the gradient of the reference feature $\nabla\langle X^{0}\rangle$. This indicates that the delay retrieval methods discussed in the preceding sections are only sensitive to the $|m|=1$ components. In the cases where feature $Y$ is in the impulsive regime, as discussed in the preceding sections, Eqn. (20) reduces to Eqn. (9) after rearranging the terms, as shown in Supplemental Sec. 4.1.
@@
To illustrate Eqn. (20), we simulate the MD for Auger-Meitner (AM) decay of a molecule in a superposition of two core-ionized states, prepared by $K$-shell ionization with a $0.2$ fs FWHM pulse [17]. The two intermediate cationic states have an energetic separation of $\Delta I_{\mathrm{P}}=7\hbar\omega_{L}$, and the AM decay process couples these two states to a common dication state. In our simulation, the dressing-free AM electrons have a central momentum of $4.34$ a.u. and $4.38$ a.u. The emitted electrons are dressed by a $|\bm{A}|=0.1$ a.u., $cT_{L}=1.85\,\mathrm{\mu m}$ dressing field (equivalent to Fig. 10) and the MD considered is the $p_{z}=0$ slice. Other details of the simulation of the AM distribution are provided in Supplemental Sec. VI. The decay of this superposition of core-ionized states results in a characteristic modulation of $Y$ over $\kappa$ at a period of $72\pi/$ 7, which results from interference between the two pathways via different intermediate states [17]. This modulation appears across a wide range of AM momenta. Here we have assumed that $Y^{0}$ is angularly isotropic. This means that the MD of AM electrons at different angular positions $\theta_{p}$ are equivalent up to an offset in $\kappa$, i.e. $Y(\theta_{p},r;\kappa,k)=Y(0,r;\kappa-\theta_{p},k)$, so the yield is also modulated over angular position $\theta_{p}$.
@@
In Fig.13 we analyze the MD by partitioning the feature $Y$ into $N_{P}=180$ angular bins, between radial boundaries $p_{\mathrm{min}}{=}4.40$ a.u. and $p_{\mathrm{max}}{=}4.50$ a.u.. Figure 13(a) shows the integral of $Y(\bm{r};\bm{k})$, in the angular bin at $\theta_{p}=0$, as a function of the streaking directions of the reference feature $\kappa$. As shown in Fig. 13(b), the yield in these angular bins $Y_{p}$ is clearly modulated with both $\kappa$ and $\theta_{p}$. We compute the streaking covariance $K[X,Y]$ between $Y$ and the impulsive reference feature $X$. This reference feature $X$ is simulated in the same dressing field ($k=0.1\,\mathrm{a.u.}\approx 0.5\Delta p_{X}$), and the ROIs $\{Q_{q}\}$ are remain between $p_{\mathrm{MG}}$ and $p_{\max}{=}2.15$ a.u. as in Fig. 10. The $72\pi/$ 7-period modulation is barely visible in $K[X,Y]$, as shown in Fig. 13(c). This attenuation of the $72\pi/$ 7-period modulation results from the low sensitivity of $K[X,Y]$ to $\mathcal{Y}_{7}$. We can study the sensitivity of $K[X,Y]$ to the components $\mathcal{Y}_{m}$ by computing $\mu_{m}$ according to Eqn. (21) and integrating over the angular bins $Q_{q}$. We show the magnitude of the resultant sensitivities in Fig. 13(d), which demonstrate a decrease with increasing $m$. Such dependence on $m$ is determined by $k,\Delta p_{X}$, and the ROIs on feature $X$, as described below. The products of multiplying these calculated $\mu_{m}$ to the corresponding Fourier components $\mathcal{Y}_{m}$ yield good agreements with the Fourier coefficients of $K[X,Y]$ with respect to $\theta_{p}$, which is visualized for the $\theta_{q}=0$ slice in panel (e), corroborating Eqn. (21).
@@
We notice that in the small streaking regime, according to Equations (20)(21), the sensitivity is given by $\int_{Q_{q}}\mu_{m}d^{2}\bm{r}\approx\frac{k^{m}}{m!2^{m/2}}\int_{Q_{q}}\partial_{+}^{m}\langle X^{0}\rangle d^{2}\bm{r}$, which is proportional to the regional integral of a $m$-th order derivative of the reference feature. Therefore, the scaling of the sensitivity over $m$ depends on the streaking amplitude relative to the characteristic momentum scale $\delta p$ far below which the variations in $\langle X^{0}\rangle$ become negligible, e.g. the width $\Delta p_{X}$ in the example shown in Fig. 13. In the limit of small streaking, the magnitude of sensitivity $|\mu_{m}|$ scales as $o((k/\delta p)^{m}/m!)$, which rapidly decreases with increasing $m$. Even in cases where $k$ is comparable to $\delta p$, as it is the case in Fig. 13, the streaking covariance tends to understate the high-order Fourier components due to the $1/m!$ factor in the sensitivity. This also indicates that to improve the sensitivity to high-order components, it is advisable to intensify the dressing field and to reduce the smoothness of the reference feature to increase $k/\delta p$.
@@
The bilinearity of covariance also leads to the independence of sensitivity $\mu_{m}$ on feature $Y$. Rather, $\mu_{m}$ is determined by the reference feature $X$ and the streaking amplitude $k$, as shown in Eqn. (21). Therefore, when detecting $Y$ as a function of $\kappa$ by analyzing the covariance with an impulsive reference feature $X$, the sensitivity to certain Fourier components $\mathcal{Y}_{m}$ can be designed in simulation, by choosing the streaking amplitude $k$, the shape of $X^{0}$ and the position of ROIs $\{Q_{q}\}$, without detailed prior knowledge in $Y$.